OpenStudy (anonymous):

i need to find the largest possible domain and range of 2x^2-2x

OpenStudy (zzr0ck3r):

is there any reason you would need to restrict the domain at all?

OpenStudy (zzr0ck3r):

and as far as range, use the vertex formula to find the lowest point, and then everything above that, including that, is the range.(because this is a porabola that opens up)

OpenStudy (anonymous):

i found the range and domain but what im having trouble with is how to find the minimum values of f(x) =2x^2-2x=-1/2 at x=1/2 if you could show me how to find these values i may be able to solve some other problems i have here